149600000 km is equivalent to:

A) 149.6 Gm

B) 1.496 Tm

C) 149.6 Mm

D) 1.496 nm


1
Expert's answer
2022-06-06T08:55:30-0400
"{149600000 \:\rm km=1.496*10^{8}\: km}\\
=1.496*10^{11}\:\rm m=149.6*10^{9}\:\rm m=149.6\: Gm"

Answer: A) 149.6 Gm